lovide an appropriate response. Agroup of students were asked if they carry a credit card. The responses are listed in the table Credit Card |Not a Credit Card Class Freshman Sophomore Total Carrier Carrier 39 18 Total 40 100 If a student is selected at random, find the probability that he or she owns a credit card given that the stude places O 0.350 0.488 0.650 O 0.210 Click Save and Submit to save and submit. Click Save All Answers to save all ansuers If a student is selected at random, find the probability that he or she owns a credit card given that the student is a freshman
Answer #1

P(own credit card | fresh man ) = P(own credit card and freshman)/P(freshman) = 21/60=0.350